OpenStudy (anonymous):

Solve the equation one-half times x plus three-fourths equals one-fourth for x. x equals one-fourth x equals negative one-fourth x equals 1 x equals negative 1

OpenStudy (texaschic101):

1/2x + 3/4 = 1/4 --- multiply the entire equation by the LCD, which is 4 (4) 1/2x + (4)3/4 = (4)1/4 2x + 3 = 1 -- subtract 3 from both sides 2x + 3 - 3 = 1 - 3 -- combine like terms 2x = -2 -- divide both sides by 2 x = -1

OpenStudy (anonymous):

oh i forgot to combine like terms :l

OpenStudy (texaschic101):

lol...you can't forget to do that

OpenStudy (anonymous):

oh.... :}